# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

### v3

@@ 3.4 Conformal Prediction and uncertainty estimation
where $|C|$ represents the cardinality of the conformal set, $N$ represents the number of classes and $\Delta_{max}$ stands for the maximum difference between the classes in the conformal set.
Algorithm 3 Conformal prediction for gradation
Input: Calibration samples $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$, error rate $\alpha$triangleright $y_{i}$ is the true grade
Step 1: Compute conformal scores:
for each calibration sample $i$ do
     $\text{score}_{i}\leftarrow 1-p_{i_{y_{i}}}$triangleright where $p_{i_{y_{i}}}$ is the predicted probability of the true class $y_{i}$
end for
Step 2: Compute adjusted quantile:
$q\leftarrow\frac{\lceil(n+1)\cdot(1-\alpha)\rceil}{n}$
$\hat{q}\leftarrow\text{quantile}(\{\text{score}_{i}\},q)$
Step 3: Form calibration sets:
for each test sample do
     if $p_{j}\geq 1-\hat{q}$ then
         Include class $j$ in the calibration set $\triangleright$ where $p_{j}$ is the predicted probability of the sample belonging to class $j$
     end if
end for
Algorithm 4 Conformal prediction for risk estimation
Input: Calibration samples $\{\mathbf{x}_{i},r_{i}\}_{i=1}^{n}$, error rate $\alpha$
$\triangleright$ $r_{i}$ is the upper bound risk of the true time bin representing maximum risk
Step 1: Compute conformal scores:
for each calibration sample $i$ do
     $\text{score}_{i}\leftarrow|r_{i}-risk_{est_{i}}|$triangleright where $risk_{est_{i}}$ is the predicted risk
end for
Step 2: Compute adjusted quantile:
$q\leftarrow\frac{\lceil(n+1)\cdot(1-\alpha)\rceil}{n}$
$\hat{q}\leftarrow\text{quantile}(\{\text{score}_{i}\},q)$
Step 3: Form calibration sets:
for each test sample do
     $risk_{lb}\leftarrow risk_{pred}-\hat{q}$
     $risk_{ub}\leftarrow risk_{pred}+\hat{q}$
     if $risk_{lb}\leq t_{lb}$ or $t_{ub}\leq risk_{ub}$ then
         Include time bin $t$ in the calibration set
$\triangleright$ where $t_{lb}$ and $t_{ub}$ are the lower bound and upper bound risks corresponding to time bin $t$
     end if
end for
3.4.1 Conformal for risk estimation
@@
Conformal prediction is performed to estimate the upper and lower bounds of the predicted risk. The marginal conformal condition for risk estimation is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(risk_{lb}\leq risk_{est}\leq risk_{ub})}$ (11)
where $risk_{est}$ is the estimated risk, and $risk_{lb}$ and $risk_{ub}$ are the lower and upper bound of risks obtained from conformal prediction. For risk estimation, the survival time in months is divided into 4 time bins. The time bins covered between the lower and upper bound of the risk give the conformal set $C$. Algorithm 4 mentions the detailed procedure for conformal risk estimation Angelopoulos and Bates [2021]. The uncertainty for risk estimation is further computed as,
$uncertainty=\frac{|C|}{N}\times\frac{|risk_{ub}-risk_{lb}|}{(-1)-(-5)}$ (12)
@@
where N=4, the number of time bins, -1 is the highest possible risk and -5 is the lowest possible risk score.
4 Discussion
Here we present PathGen, a crossmodal generative diffusion model for synthesizing transcriptomic features from WSIs of H&\&&E stained slides. We demonstrated that PathGen synthesized transcriptomic features are highly correlated to true patient transcriptomes. By using a multimodal AI model to predict cancer grade and survival risk from WSIs and transcriptomic features, with state-of-the-art performance, we demonstrated that combining the features learnt from WSIs using a vision transformer based foundation model with the synthesized transcriptomic features from PathGen, significantly improved predictive accuracy compared to WSIs alone. Remarkably, we found that there was no significant difference in accuracy when real patient transcriptomic data were used instead of PathGen synthesized transcriptomic data in the multimodal model, demonstrating that PathGen captures the independent gradation/prognostic information in true patient gene expressions.
H&\&&E stained slides of tumour tissue are essential for diagnostic and prognostic assessment of malignancy, and are collected routinely. However, transcriptomic tests, though they have the potential to inform prognosis and guide treatment selection Sparano et al. [2018], Tsakiroglou et al. [2023] are not routinely performed, especially in the public sector Harnan et al. [2019]. This is due to a number of reasons, including economic cost, infrastructural requirements and incomplete knowledge about patient benefit. PathGen can synthesize realistic and informative patient transcriptomic features, from routine and inexpensive WSIs of H&\&&E stained slides. Thus in the future, we envision that PathGen can be studied as a low cost screening tool, to guide the targeted collection of transcriptomic data modalities, by potentially selecting patients who would benefit from transcriptomic assessment for more accurate diagnosis/prognosis. This would of course need a dedicated follow up study which would include patient response to treatment when selected in this manner, that would entail a separate clinical validation project.
For PathGen to be useful in clinic, it must be trusted. To facilitate this trust we have undertaken two measures. Firstly, by employing co-attention tools for model transparency, we map gradation and survival risk estimates back onto the WSI images in a distributed manner. This has provided a window into intra-tumour heterogeneity, and can highlight cases where whole slide assessment and distributed assessment of tumour grade may differ. Secondly, we have employed conformal prediction for assessment of model reliability, providing not simply point estimates for gradation and survival risk, but ranges which contain true values with a specified certainty. This analysis has also allowed us to assess model fairness, by comparing model reliability across patient demographics such as age and gender. Importantly, we found that there was no significant difference in model reliability or fairness when using PathGen synthesized transcriptomic features compared to real patient transcriptomes. For reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository at https://github.com/Samiran-Dey/PathGen.
Our initial results are based on assessment of two publicly available cancer datasets from TCGA, describing glial and renal tumours. Though it is encouraging that PathGen shows state-of-the-art performance across diverse tumour types, further assessment in other malignancies and in independent datasets, would expand the remit of our results. PathGen is able to reliably synthesize transcriptomic embeddings, describing 6 distinct gene sets. Though these gene sets (both synthesized and real) contain important diagnostic/prognostic information, they do not capture the full diversity of the transcriptome. There is likely a limit to how much of the full transcriptome can be synthesized from only knowledge of WSIs, and hence we do not claim that PathGen replaces the need for real transcriptomic data, but rather it provides a potential way of estimating the benefit of collecting the full data across the population. The increasing use of techniques such as spatial transcriptomics Williams et al. [2022] will provide insight into this question, as well as WSI patch level resolution of transcriptomics, to guide development of synthesized transcriptomic data with a greater number of features in the near future.
Acknowledgment
@@
The authors declare no conflict of interest. S Dey is funded by the IIT KGP AI4ICPS Chanakya Fellowship. P Basuchowdhuri is supported by the Indo-Swedish DBT-Vinnova project, BT/PR41025/Swdn/135/9/2020, for supporting this research. CRSB was supported by the CRUK City of London Centre Award [CTRQQR-2021/100004]. T Chakraborti is supported by the Turing-Roche Strategic Partnership. We are grateful to Prof Benjamin MacArthur from the Alan Turing Institute and Mr. Chris Harbron from Roche Pharmaceuticals for their insightful comments and constructive feedback throughout this project.
@@
Table 1: Table of results for evaluation of the use of synthesized transcriptomic data for grade and survival prediction in four different setting 1) We do not use transcriptomic data at all and attend to the WSI patches randomly; 2) We use the gene expression levels only to co-attend to the WSI patch embeddings; 3) We use the transcriptomic embeddings collectively with co-attended WSI features for survival risk estimation only; 4) We use transcriptomic features for co-attending to WSI patches, and then use the transcriptomic embeddings along with the co-attended features for performing both survival risk estimation and gradation. Setting 3 gives the best mean AUC and C Index for gradation and survival risk estimation across both data cohorts. Further, ANOSIM test for combined risk estimation and gradation between the two cases suggested that the difference in performance between all the cases is significant (p-value $<0.05$). Thus, the architecture of our model is aligned to setting 3. ↑ indicates the larger the better. Best mean scores are marked in bold.
@@
Figure 8: Ablation for $\lambda$. $\lambda=0$ and $\lambda=1$ correspond to the extreme cases of equation $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$ where the model is trained only using gradation loss and only using survival loss respectively. Using only gradation loss leads to poor performance of risk estimation when $\lambda=0$ and using only risk estimation loss leads to a poor gradation performance. Learning both grade and survival risk together helps the model to correlate between the two and learn better. This is expected as one of the losses essentially becomes unused, so we decide to ignore the extreme cases of $\lambda=0$ and $\lambda=1$. Also, we observe that for $\lambda=0.3$ the model reaches an optimal performance where both AUC and C Index scores are close and high together for both cohorts.
Table 2: Table of results for comparison of synthesized and real transcriptomic features. The correlation coefficient is calculated using Spearman correlation. The p-values are evaluated by performing an unpaired T-test for the aggregate gene expression level of each group. ↑ indicates the larger the better and ↓ indicates the smaller the better. A higher correlation with a high corresponding MAE would denote that the range of gene expression levels is large for the particular group.
Dataset TCGA - GBMLGG TCGA - KIRC
Gene type MAE ↓ Correlation ↑ p-value ↓ MAE ↓ Correlation ↑ p-value ↓
Tumor Suppressor Genes 0.137 0.472 4.32e-06 0.116 0.629 9.21e-69
Oncogenes 0.085 0.770 8.05e-53 0.074 0.817 7.08e-54
Protein Kinases 0.084 0.749 4.83e-79 0.078 0.785 6.48e-84
Cell Differentiation Markers 0.069 0.724 8.55e-50 0.070 0.777 1.38e-53
Transcription Factors 0.122 0.615 4.52e-47 0.117 0.666 6.107e-52
Cytokines and Growth Factors 0.062 0.771 2.43e-66 0.095 0.610 2.48e-72
All genes 0.098 0.713 1.18e-67 0.097 0.717 3.65e-69
Figure 9: Predicted intra-tumour heterogeneity at different magnification levels (a) A whole slide image (WSI) of ground truth grade II, survival time 62.62 months and alive survival status belonging to a male patient of 47 years. (b) A WSI of ground truth grade IV, survival time 54.61 months and dead survival status belonging to a 55 years old female patient. The WSIs belong to TCGA-KIRC. It is observed that for higher resolution images the distributed predictions are more accurate.
Figure 10: Explainability analysis on TCGA-GBMLGG data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomes and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
Figure 11: Explainability analysis on TCGA-KIRC data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomes and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
@@
Table 3: Table of results for co-attention maps generated while using synthesized and real transcriptomic features. The correlation coefficient is calculated using Spearman correlation. ↑ indicates the larger the better and ↓ indicates the smaller the better. A higher correlation with a high corresponding MAE would denote that the range of gene expression levels is large for the particular group.
Dataset TCGA - GBMLGG TCGA - KIRC
Gene type MAE ↓ Correlation ↑ MAE ↓ Correlation ↑
Tumor Suppressor Genes 0.801 0.998 0.087 0.946
Oncogenes 2.302 0.999 0.460 0.995
Protein Kinases 2.526 0.894 0.429 0.881
Cell Differentiation Markers 2.305 0.969 0.319 0.610
Transcription Factors 1.973 0.997 1.865 0.990
Cytokines and Growth Factors 2.505 0.531 0.387 0.404
All genes 2.069 0.939 0.591 0.851
Table 4: Table of results across various groups for grade prediction. p-values are computed using the Wilcoxon rank-sum test to compare the scores for grades predicted using real and synthesized transcriptomic data. ↑ indicates the larger the better and ↓ indicates the smaller the better. Equivalent, better or close ($<0.01$) performance with synthesized data is marked in bold.

### v4

@@ METHODS
where $|C|$ represents the cardinality of the conformal set, $N$ represents the number of grade classes, and $\Delta_{max}$ stands for the maximum difference between the grades in the conformal set.
[…]
Conformal for survival risk estimation
@@
For survival risk estimation, conformal prediction is performed to estimate the upper and lower bounds for the estimated risk. The marginal coverage property for risk estimation ensures that the ground truth survival is covered within the predicted bounds by a probability of more than $1-\alpha$, where $\alpha$ is a chosen error rate, and is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(Y_{test}\in\;[risk_{lb},risk_{ub}])}$ (12)
where $risk_{lb}$ and $risk_{ub}$ are the conformalized lower and upper bounds of the risk predicted from $X_{test}$. For risk estimation, the survival time in months is divided into 4 time bins for a given population, considering the vital status. The time bins covered between the lower and upper bounds of the risk give the conformal set $C$. Similar to gradation, stratified conditional conformal prediction is performed to ensure fairness in coverage over different demographic groups. The stratified conditional coverage property for survival risk estimation is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(Y_{test}\in[risk_{lb},risk_{ub}]\;|\;X_{test}\in g_{i}}),\quad\forall\,g_{i}\in G$ (13)
where G represents the demographic category and $g_{i}$ represents the groups in the category. Box Conformal for gradation mentions the detailed conformal procedure for survival risk estimation. The uncertainty for risk estimation is further computed as,
$uncertainty=\frac{|C|}{N}\times\frac{|risk_{ub}-risk_{lb}|}{|risk_{high}-risk_{low}|}$ (14)
@@
where N=4, the number of time bins, $risk_{high}=$ is the highest possible risk, and $risk_{low}=$ is the lowest possible risk score.
Both the uncertainty measures given in equation 11 and equation 14 are metrics to quantify uncertainty for gradation and survival risk estimation, respectively.
Estimation of coverage slack
However, though conformal properties (equations 9, 10, 12 and 13) ensure that the obtained coverage is at least $1-\alpha$, the coverage of conformal prediction conditionally on the calibration set is a random quantity Angelopoulos and Bates [2021]. The deviation in coverage is a function of the cardinality of the calibration set (n), and the error rate ($\alpha$), and follows the distribution: $\mathrm{Beta}(n+1-l,\;l\;)$, where $l=\lfloor(n+1)\alpha\rfloor$ Vovk [2012]. The expected deviation ($\varepsilon$) can be computed using the equation,
$\varepsilon(n,\alpha)=\max((1-\alpha)-L,\;U-(1-\alpha),\;0)$ (15)
where L and U are the lower and upper bounds for the 90% confidence-type interval around $1-\alpha$, for $\alpha=0.1$, and is given by,
$L,U]=[F^{-1}_{\mathrm{Beta}(a,b)}(0.05),\;F^{-1}_{\mathrm{Beta}(a,b)}(0.95)]$ (16)
Hence, conformalized coverage slack of $\alpha\;\pm\;\varepsilon$ is inherent for finite set calibration. The expected deviations in coverage have been computed for different calibration set sizes, using equation 15, and presented in Supplementary Table LABEL:s-table_deviation, for $\alpha=0.1$.
INCLUSION AND ETHICS
The author team includes researchers from the UK and India, and contributions reflect international collaboration. Although the data sets analysed were publicly available, we acknowledge that patient cohorts may not fully represent global populations.
DATA AVAILABILITY
For all our experiments, we use data from two anonymised publicly available open-access datasets - The Cancer Genomic Atlas (TCGA) and the Clinical Proteomic Tumor Analysis Consortium (CPTAC). The TCGA data can be accessed at https://www.cancer.gov/tcga, and the CPTAC data can be accessed at https://gdc.cancer.gov/about-gdc/contributed-genomic-data-cancer-research/clinical-proteomic-tumor-analysis-consortium-cptac. We use the corresponding transcriptomic data downloaded from cBioPortal (https://www.cbioportal.org) for all cohorts and datasets. The gene categories used are found at https://github.com/mahmoodlab/MCAT/blob/master/datasets_csv_sig/signatures.csv. No primary data was collected as part of this study. Source data are provided with this paper.
CODE AVAILABILITY
For the reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository at https://github.com/Samiran-Dey/PathGen Dey et al. [2025] under the CC-BY-NC-ND 4.0 license.
ACKNOWLEDGEMENTS
@@
S Dey was supported by the IIT KGP AI4ICPS Chanakya Fellowship. P Basuchowdhuri is supported by the Indo-Swedish DBT-Vinnova project, BT/PR41025/Swdn/135/9/2020. CRSB was supported by the CRUK City of London Centre Award [CTRQQR-2021/100004]. T Chakraborti is supported by the Turing-Roche Strategic Partnership. We are grateful to Prof Benjamin MacArthur from the Alan Turing Institute and Mr. Chris Harbron from Roche Pharmaceuticals for their insightful comments and constructive feedback throughout this project.
AUTHOR CONTRIBUTIONS
S. Dey has contributed to conceptualization, methodology development, experimental studies, and writing. P. Basuchowdhuri, S.K. Saha, D. Parashar contributed to overall feedback, knowledge sharing and writing/editing of the manuscript. T. Chakraborti has contributed to conceptualization, problem setting, overall research direction, project supervision and writing/editing the manuscript. C.R.S. Baernji has provided clinical expertise and feedback, and contributed to writing/editing the manuscript.
COMPETING INTERESTS
The authors declare no competing interests.
FIGURES
Figure 1: Methodology overview. Pipeline for the proposed methodology. Our diffusion-based crossmodal generative model, PathGen, synthesizes transcriptomic features from whole slide image patch embeddings obtained using a state-of-the-art foundation model from the whole slide images. Multimodal prediction is performed for both diagnosis (cancer grading) and prognosis (survival risk) using the synthesized transcriptomic data and histopathology images to predict the added value of transcriptomic data, for each patient. Uncertainty quantification provides a patient-level estimate of model reliability and provides clinicians with survival risk bounds and tumour grade prediction sets which are guaranteed to contain the ground truth with a specified probability. Distributed predictions also provide a window to understand the intra-tumour heterogeneity.
Figure 2: Evaluation of synthesized transcriptomic data. (a) Comparison between real and synthesized gene expression levels over all genes and for different gene groups. (b) Plot of significant performance improvement on using synthesized transcriptomic data over using WSIs alone (marked as without), and closeness of using synthesized and real transcriptomic data for survival risk estimation and gradation. (c) Plot for comparison of uncertainty and obtained coverage for survival risk and grade predictions using real and synthesized transcriptomic data. Source data are provided as a Source Data file.
Figure 3: Explainability maps for a male patient of age 58, ground truth grade III, survival time 10.71 months, and survival status alive from the TCGA-LGG cohort. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
Figure 4: Explainability maps for a male patient of age 40, ground truth grade IV, survival time 33.99 months, and survival status deceased from the TCGA-KIRC cohort. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
Figure 5: Explainability analysis. (a) (i) Plot for comparison between distributed and non-distributed WSI predictions for gradation using both real and synthesized transcriptomic data. (ii) Plot showing normalized average co-attention value for the true grade is higher than that of false grades. (b) Comparison between co-attention maps obtained using real and synthesized transcriptomic data over all genes and for different gene groups. (c) Plot illustrating the percentage contribution of the gene groups in co-attention for different TCGA and CPTAC data cohorts. Source data are provided as a Source Data file.
Figure 6: Illustrative result for evaluating fairness in model predictions across demographic categories and data cohorts. The cases have been arranged in ascending order of predicted patient risk. The estimated uncertainty in survival risk and grade prediction is also shown. Source data are provided as a Source Data file.
Figure 7: Model Architecture. (a) Architecture of our diffusion model, PathGen, used for synthesizing gene expression levels from histopathology images. (b) Architecture of the modified MCAT model for gradation and risk estimation, MCAT_GR. (c) Detailed architecture of PathGen transformer. (d) Architecture of gene encoder that converts gene expression level to gene embeddings. (e) Representation of genomic-guided co-attention function. (f) Architecture of gene decoder that converts gene embeddings back to gene expression levels.
Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions
@@
Table 1: Details of data splits across datasets (TCGA, CPTAC) and cancer cohorts used in the experiments. The cohorts include glioblastoma-glioma (TCGA-GBMLGG, CPTAC-GBM), renal (TCGA-KIRC), uterine (TCGA-UCEC, CPTAC-UCEC) and breast (TCGA-BRCA). Data from the CPTAC cohorts were excluded from training. Models trained on the TCGA cohorts were evaluated on the corresponding CPTAC datasets.
Datasets Cases Slides
Training Validation Calibration Testing Training Validation Calibration Testing
TCGA-GBMLGG 532 56 75 82 653 69 86 104
TCGA-KIRC 337 34 45 46 344 42 48 51
TCGA-UCEC 191 19 30 27 210 23 31 30
TCGA-BRCA 676 78 97 95 729 80 102 99
CPTAC-GBM - - 32 30 - - 151 91
CPTAC-UCEC - - 38 33 - - 194 170
Table 2: Table for number of genes in each gene group in the transcriptomic data. Same number of genes have been considered in each gene groups for same cohorts from different datasets. We note that there is repetition of genes across gene groups. The last row shows the additive result of the number of genes across gene groups including repetitions, and not the total number of genes considered for prediction.
Gene type TCGA-GBMLGG TCGA-KIRC TCGA-UCEC TCGA-BRCA CPTAC-GBM CPTAC-UCEC
Tumor Suppressor Genes 84 145 152 227 84 152
Oncogenes 314 522 554 959 314 554
Protein Kinases 498 669 700 934 498 700
Cell Differentiation Markers 424 681 693 843 424 693
Transcription Factors 1396 2366 2489 3712 1396 2489
Cytokines and Growth Factors 428 990 965 1142 428 965
All gene groups 3144 5373 5553 7817 3144 5553
Table 3: Demographic details of the data used for calibration and testing for all the datasets and cohorts. Ground truth grades are absent for TCGA-BRCA and CPTAC-GBM datasets. Also, patient age is missing in the clinical data for the CPTAC cohorts. ncal denotes the number of samples in the calibration set and ntest denotes the number of samples in the test set.
Category Group TCGA - GBMLGG TCGA - KIRC TCGA - UCEC TCGA - BRCA CPTAC - GBM CPTAC - UCEC
ncal ntest ncal ntest ncal ntest ncal ntest ncal ntest ncal ntest
overall - 86 104 144 153 93 90 306 297 453 273 582 510
gender male 55 66 102 99 - - 282 120 282 120 - -
female 31 38 42 54 93 90 171 153 171 153 582 510
age <40 33 43 6 3 - - 30 9 - - - -
40 to 60 28 37 51 84 42 39 162 144 - - - -
>60 25 24 87 66 51 51 114 144 - - - -
magnification level 1 27 29 48 51 31 30 102 99 151 91 194 170
level 2 27 38 48 51 31 30 102 99 151 91 194 170
level 3 32 37 48 51 31 30 102 99 151 91 194 170
censorship alive 48 67 75 102 81 78 279 279 165 69 555 447
dead 38 37 69 51 12 12 27 18 288 204 27 63
grade grade I - - 3 3 30 24 - - - - 225 99
grade II 34 39 57 63 21 24 - - - - 207 270
grade III 28 36 60 63 42 42 - - - - 150 141
grade IV 24 29 24 24 - - - - 453 273 - -
time bin bin 1 80 82 72 72 66 63 255 249 453 273 474 423
bin 2 4 13 39 60 27 27 45 45 - - 108 87
bin 3 2 9 21 18 - - 3 3 - - - -
bin 4 - - 12 3 - - - - - - - -
Table 4: Table of results for evaluation of the use of synthesized transcriptomic data for grade and survival prediction in four different settings- 1) We do not use transcriptomic data at all and attend to the WSI patches randomly; 2) We use the gene expression levels only to co-attend to the WSI patch embeddings; 3) We use the transcriptomic embeddings collectively with co-attended WSI features for survival risk estimation only; 4) We use transcriptomic features for co-attending to WSI patches, and then use the transcriptomic embeddings along with the co-attended features for performing both survival risk estimation and gradation. Setting 3 gives the best mean AUC and C Index for gradation and survival risk estimation across both data cohorts. Further, ANOSIM test for combined risk estimation and gradation between the two cases suggested that the difference in performance between all the cases is significant (p-value <0.05<0.05). Thus, the architecture of our model is aligned to setting 3. ↑ indicates the larger the better. Best mean scores are marked in bold.
@@
Figure 8: Ablation for $\lambda$. Plot for AUC and C Index across $\lambda$ for (a) TCGA-GBMLGG, (b) TCGA-KIRC. $\lambda=0$ and $\lambda=1$ correspond to the extreme cases of equation $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$ where the model is trained only using gradation loss and only using survival loss respectively. Using only gradation loss leads to poor performance of risk estimation when $\lambda=0$, and using only risk estimation loss leads to poor gradation performance. Learning both grade and survival risk together helps the model correlate between the two and learn better. This is expected as one of the losses essentially becomes unused, so we decide to ignore the extreme cases of $\lambda=0$ and $\lambda=1$. Also, we observe that for $\lambda=0.3$ the model reaches an optimal performance where both AUC and C Index scores are close and high together for both cohorts.
@@
Table 5: Table of results for comparison of synthesized and real transcriptomic features. Normalized MAE results are denoted with nMAE. The correlation coefficient is calculated using Spearman correlation, and all corresponding correlation p-values are found to be <0.05<0.05, denoting significant correlation. ↑ indicates the larger the better, and ↓ indicates the smaller the better. A higher correlation with a high corresponding MAE would denote that the range of gene expression levels is large for the particular group.
Gene type TCGA-GBMLGG TCGA-KIRC TCGA-UCEC
nMAE Correlation nMAE Correlation nMAE Correlation
Tumor Suppressor Genes 0.184 0.472 0.166 0.629 0.177 0.372
Oncogenes 0.160 0.770 0.136 0.817 0.182 0.397
Protein Kinases 0.148 0.749 0.183 0.785 0.183 0.306
Cell Differentiation Markers 0.137 0.724 0.130 0.777 0.194 0.339
Transcription Factors 0.143 0.615 0.197 0.666 0.139 0.623
Cytokines and Growth Factors 0.107 0.771 0.091 0.610 0.236 0.270
All gene groups 0.141 0.713 0.160 0.717 0.173 0.436
Gene type TCGA-BRCA CPTAC-GBM CPTAC-UCEC
nMAE Correlation nMAE Correlation nMAE Correlation
Tumor Suppressor Genes 0.194 0.568 0.259 0.445 0.228 0.540
Oncogenes 0.136 0.705 0.043 0.874 0.197 0.582
Protein Kinases 0.131 0.766 0.036 0.923 0.209 0.516
Cell Differentiation Markers 0.166 0.617 0.036 0.941 0.206 0.541
Transcription Factors 0.167 0.629 0.316 0.490 0.146 0.827
Cytokines and Growth Factors 0.136 0.628 0.035 0.965 0.197 0.493
All gene groups 0.155 0.642 0.167 0.662 0.178 0.669
Table 6: Table of results for comparing predictive performance on using only WSIs (without transcriptomic data), using synthesized transcriptomic data and WSIs, and using real transcriptomic data and WSIs. The separate p-values for gradation and risk estimation are computed by performing two-sided Wilcoxon rank-sum test, whereas the p-values over gradation and risk estimation together are computed using two-sided ANOSIM test. ANOSIM test $p$-values =0.01=0.01, for comparison between without vs. synthesized, for TCGA-GBMLGG, TCGA-KIRC, TCGA-UCEC, CPTAC-GBM, and CPTAC-UCEC. However, for real vs. synthesized, ANOSIM test $p$-values evaluate to 0.929 for TCGA-GBMLGG, 0.327 for TCGA-KIRC, 1.000 for TCGA-UCEC, 0.993 for CPTAC-GBM and 1.000 for CPTAC-UCEC. ↑indicates the larger the better, and ↓indicates the smaller the better.
Gradation
AUC↑ Wilcoxon test p-value
Datasets without synthesized real without vs. synthesized ↓ real vs. synthesized ↑
TCGA-GBMLGG 0.823 0.890 0.907 4.23E-02 9.95E-01
TCGA-KIRC 0.714 0.773 0.778 2.21E-04 8.50E-01
TCGA-UCEC 0.796 0.821 0.828 1.28E-02 9.41E-01
CPTAC-GBM 0.736 0.865 0.863 8.40E-03 9.85E-01
CPTAC-UCEC 0.508 0.593 0.593 7.76E-42 1.00E+00
Risk Estimation
C Index ↑ Wilcoxon test p-value
Datasets without synthesized real without vs. synthesized ↓ real vs. synthesized ↑
TCGA-GBMLGG 0.842 0.861 0.866 8.20E-07 6.04E-01
TCGA-KIRC 0.671 0.681 0.697 1.23E-11 3.00E-01
TCGA-UCEC 0.663 0.673 0.680 4.81E-31 8.87E-01
TCGA-BRCA 0.603 0.720 0.720 2.35E-03 9.30E-01
CPTAC-GBM 0.547 0.565 0.564 3.90E-33 8.15E-01
CPTAC-UCEC 0.518 0.530 0.533 2.26E-164 9.32E-01
Table 7: Table for results for gradation and survival risk estimation using real and synthesized transcriptomic data, separately for the TCGA-GBMLGG cohorts. The $p$-values are computed using two-sided Wilcoxon rank-sum test.
TCGA-GBMLGG Gradation (AUC↑) Risk Estimation (C Index ↑)
cohorts real synthesized Wilcoxon p-value real synthesized Wilcoxon p-value
TCGA-LGG 0.835 0.81 0.775 0.707 0.675 0.472
TCGA-GBM 0.995 0.993 0.822 0.509 0.539 0.489
Table 8: Table for ablation study on percentage contribution of the different gene groups in co-attention with WSI features for different cohorts.
Gene type Percentage Contribution (in %)
TCGA-GBMLGG TCGA-KIRC TCGA-UCEC TCGA-BRCA CPTAC-GBM CPTAC-UCEC
Tumor Suppressor Genes 12.51 6.05 5.25 11.56 12.15 4.74
Oncogenes 18.02 15.59 4.33 12.90 13.16 6.35
Protein Kinases 10.40 11.86 38.41 21.12 15.30 29.09
Cell Differentiation Markers 10.00 8.10 10.82 18.14 23.55 18.00
Transcription Factors 42.21 48.80 17.74 23.49 29.06 25.89
Cytokines and Growth Factors 6.85 9.61 23.44 12.80 6.77 15.93
Figure 9: Predicted intra-tumour heterogeneity at different magnification levels (a) A whole slide image (WSI) of ground truth grade II, survival time 62.62 months and alive survival status belonging to a male patient of 47 years. (b) A WSI of ground truth grade IV, survival time 54.61 months and dead survival status belonging to a 55 years old female patient. The WSIs belong to TCGA-KIRC. It is observed that for higher resolution images, the distributed predictions are more accurate.
Figure 10: Explainability maps for a male patient of age 62, ground truth grade IV, survival time 10.28 months, and survival status deceased from the TCGA-GBM cohort. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
Figure 11: Explainability maps for a female patient of age 50, ground truth grade II, survival time 22.78 months, and survival status alive from the TCGA-KIRC cohort. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
Figure 12: Explainability maps for a female patient of ground truth grade III, survival time 58.08 months, and survival status alive from the CPTAC-UCEC cohort. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomes for chosen regions marked with different coloured rectangles.
[…]
Category Group Gradation